OpenStudy (anonymous):

(x^2/y^6) divided by (x^7/y^3)

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ y ^{6} }\div \frac{ x ^{7}}{ y ^{3} }\] what is the general rule for dividing fractions?

OpenStudy (anonymous):

@itsallycatt

OpenStudy (anonymous):

common denominator right?

OpenStudy (anonymous):

that is addition or subtraction for division you flip the second fraction and multipy

OpenStudy (anonymous):

Oh okayy

OpenStudy (anonymous):

so if you flip the 2nd fraction it looks like this right \[\frac{ x ^{2}}{ y ^{6} }\times \frac{ y ^{3} }{ x ^{7} }\] now can anything cancel out

OpenStudy (anonymous):

\[(x^{2}/y ^{6})/(x ^{7}/y ^{3})=(x ^{2}/y ^{6})*(y ^{3}/x ^{7}) = 1/(x ^{5})(y ^{2})\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're Welcome :) @itsallycatt